Find the base of a triangle when the area given is 345 centimeters square. The Heights is 50 centimeters
Answer by ewatrrr(24785)

Triangle
Area = 1/2 of the base X the height
A = 1/2 bh
345 = (1/2)50b (In cm)
345cm^2/25cm =b
13.8cm = base
CHECKING
(1/2)50cm*13.8cm = 345cm^2 CHECKS!